OpenStudy (megannicole51):

what is an equation for the translation of (x-2)^2+(y+1)^2=16 by 3 units left and 5 units up

OpenStudy (reemii):

the answer is of the type: \((x-2 - a)^2 + (y+1 -b)^2 = 16.\) The job is to find \(a\) and \(b\).

OpenStudy (megannicole51):

okay:) so how do i do that?

OpenStudy (reemii):

I (almost) sure, when you have * (x-a) (with a positive a) the thing moved to the right. * (x+a) (with positive a) the thing moved to the left. * (y-b) it's up * (y+b) it's down.

OpenStudy (reemii):

move it to the left: write +3 in the first parenthesis. similar thing in the second parenth.

OpenStudy (reemii):

(The same thing works for translating the graph of a function. f(x-a) : shift of a to the right)

OpenStudy (megannicole51):

so it would be (x+1)^2 + (y-4)^2=16

OpenStudy (reemii):

what's your answer equation?

OpenStudy (reemii):

@megannicole51

OpenStudy (megannicole51):

like what (a) (b) (c) and (d) are?

OpenStudy (reemii):

oh i didn't see your answer: "(x+1)^2 + (y-4)^2=16" yeah that's right. what is (a) (b) (c) (d) ?

OpenStudy (megannicole51):

the answer choices lol sorry i couldnt think of the word. and sweet:) thanks!

OpenStudy (reemii):

if u have another one it'll be my last..